### v3

# Efficient Defection: Overage-Proportional Rationing Attains the Cooperative Frontier

Abstract
We study a noncooperative $n$-player game of slack allocation in which each player $j$ has entitlement $L_{j}>0$ and chooses a claim $C_{j}\geq 0$. Let $v_{j}=(C_{j}-L_{j})_{+}$ (overage) and $s_{j}=(L_{j}-C_{j})_{+}$ (slack); set $X=\sum_{j}v_{j}$ and $I=\sum_{j}s_{j}$. At the end of the period an overage-proportional clearing rule allocates cooperative surplus $I$ to defectors in proportion to $v_{j}$; cooperators receive $C_{j}$. We show: (i) the selfish outcome reproduces the cooperative payoff vector $(L_{1},\dots,L_{n})$; (ii) with bounded actions, defection is a weakly dominant strategy; (iii) within the $\alpha$-power family, the linear rule ($\alpha=1$) is the unique boundary-continuous member; and (iv) the dominant-strategy outcome is Strong Nash under transferable utility and hence coalition-proof [1]. We give a policy interpretation for carbon rationing with a penalty collar.
1 Introduction
[…]
Our main result is that this proportional slack clearing implements the cooperative frontier in dominant strategies (under bounded actions): each player’s payoff equals their entitlement in equilibrium, even though the behavior is self-regarding defection. We show the dominant-strategy profile is robust to coalition deviations under transferable utility (coalition-proof in the sense of Bernheim et al. [1]). We also characterize proportionality within a natural $\alpha$-power family: continuity at the $X=I$ boundary uniquely selects the linear rule $\alpha=1$ (Theorem 4).
We assume credible end-of-period enforcement of the clearing rule and observable claims/emissions. Dominance requires bounded actions $C_{j}\in[0,M]$; without bounds, best replies may exist only in the limit (Appendix C). Coalition-proofness is stated at the dominant-strategy profile under transferable utility.
Contributions.
(i) Implementation by efficient defection. With bounded actions, the max-claim action is a weakly dominant strategy; the induced outcome reproduces the cooperative payoff vector $(L_{j})_{j}$ and is budget balanced when $X\geq I$. (ii) Robustness to collusion. At the dominant-strategy profile, no coalition can Pareto-improve under TU; the profile is Strong Nash and hence Coalition-Proof [1]. (iii) Characterization. We consider a generalized $\alpha$-power family where the surplus $I$ is allocated proportionally to the $\alpha$-power of individual overages (i.e., $(v_{j})^{\alpha}$). We show that the linear rule ($\alpha=1$) is the unique member of this family that is continuous at the boundary $X=I$. (iv) Policy reading. As an end-of-period clearing mechanism with a penalty band, the design is compatible with forward trading and eliminates “wait-and-emit” arbitrage (Appendix D).
Relation to existing work.
The paper intersects three literatures. First, in the claims/rationing tradition (bankruptcy and uniform rationing), proportional rules are classically justified by axioms such as anonymity, consistency, and resource monotonicity [see, e.g., 2, 3, 4, 5, 6]. Our mechanism is noncooperative, budget balanced under scarcity, treats cooperators lexicographically (no-sucker-loss), and yields a new characterization via boundary continuity. Second, in congestion/CPR and network allocation, proportional sharing appears via prices and progressive filling [e.g., 7, 8], but agent payoffs there are typically price-mediated and not dominance-implementable. Our rule is price-free, direct, and dominance-implementable under bounds. Third, on coalition-proofness, we work within the Bernheim et al. [1] framework and show the DS outcome is Strong Nash under TU because coalition surplus “leaks” to nonmembers via proportional coverage.
Related literature
Claims, bankruptcy, and rationing. Classical bankruptcy/claims problems allocate a fixed estate to claimants under axioms such as anonymity, consistency, and resource monotonicity; proportional and related rules are characterized in this tradition [4, 5, 6, 3, 2]. Our setting differs: actions are strategic, cooperators are guaranteed their claims (no-sucker-loss), and budget balance holds only when scarcity binds; within this design, boundary continuity selects proportionality.
Congestion/CPR and networks. Proportional sharing appears in congestion control and progressive-filling allocations [e.g., 7, 8]; those models rely on prices and potential-game structures. We instead give a direct, price-free mechanism with dominance under bounds and coalition-proofness at equilibrium.
Coalition-proofness. We adopt the coalition-proof Nash framework of Bernheim et al. [1] and show the dominant-strategy outcome is Strong Nash under TU, hence coalition-proof, because coalition-generated surplus is diluted proportionally to overage, limiting the coalition’s net gain.
Roadmap.
Section 2 defines the rule and states the budget identity. Section 3 gives the main properties (dominance under bounds, coalition-proofness, boundary characterization). Appendix A develops the $\alpha$-family and the continuity uniqueness; Appendix B proves coalition-proofness; Appendix C provides the bounded-action regularization; Appendix D gives the policy economics of the penalty band.
2 Mechanism (slack allocation)
For each player $j$, define
$v_{j}:=(C_{j}-L_{j})_{+},\qquad s_{j}:=(L_{j}-C_{j})_{+},\qquad X:=\sum_{m=1}^{n}v_{m},\qquad I:=\sum_{m=1}^{n}s_{m}.$
Here $(x)_{+}:=\max\{x,0\}$.
Define the cooperator and defector sets by $S:=\{j:\,C_{j}\leq L_{j}\}$ and $D:=\{j:\,C_{j}>L_{j}\}$.
Assumption (costless claims). Settlement payoffs are $\pi_{j}$; submitting a claim $C_{j}$ carries no magnitude-dependent cost. (If utility were $U_{j}=\pi_{j}-\epsilon C_{j}$ with $\epsilon>0$, maximal claiming $C_{j}=M$ would not be weakly dominant.)
Bounded actions.
When we study dominant-strategy behavior we restrict claims to $C_{j}\in[0,M]$ with a common bound $M>\max_{j}L_{j}$; this is the only place where boundedness is used (see Appendix C).
For $\alpha=1$ (linear rule), defectors receive
$\widehat{v}_{j}=\begin{cases}v_{j},&X\leq I,\\[3.0pt] \dfrac{I}{X}\,v_{j},&X>I,\end{cases}\qquad\text{and}\qquad\pi_{j}=\begin{cases}C_{j},&C_{j}\leq L_{j},\\[2.0pt] L_{j}+\widehat{v}_{j},&C_{j}>L_{j}.\end{cases}$ (1)
Aggregate payoffs satisfy
$\sum_{j}\pi_{j}\;=\;\sum_{j}L_{j}\;-\;\max\{I-X,0\},$ (2)
If $X<I$, the gap IXI-X is unused surplus; if $X\geq I$ (scarcity binds), the rule is budget balanced. At the cooperative profile $C=L$ we have $(X,I)=(0,0)$ and $\pi_{j}=L_{j}$; at the all-defect profile ($I=0$) we again have $\pi_{j}=L_{j}$.
Design trade-off (incentives vs ex-post efficiency).
The mechanism attains the cooperative frontier in equilibrium by tolerating off-equilibrium inefficiency: when $X<I$, the gap IXI-X is discarded rather than rebated. This potential waste creates strong ex-ante incentives to claim aggressively; in the dominant-strategy outcome all agents claim $M$, yet realized payoffs equal $L_{j}$ and total welfare $\sum_{j}L_{j}$ is achieved. The result is an incentives-for-efficiency trade-off, not a free lunch.
Normative rationale for continuity at $X=I$.
Boundary continuity eliminates settlement cliffs under measurement/reporting noise. With small symmetric noise near $X=I$, any $\alpha\neq 1$ creates a boundary jump that yields a finite expected coverage bias even for small noise, whereas the linear rule ($\alpha=1$) removes the jump so the bias vanishes with the noise and incentives are locally robust.
Axiomatic contrast to CPR/congestion.
(i) No-Sucker-Loss: if $C_{j}\!\leq\!L_{j}$ then $\pi_{j}=C_{j}$ regardless of others. (ii) Scarcity-Budget-Balance: if $X\!\geq\!I$ then $\sum_{j}\pi_{j}=\sum_{j}L_{j}$. Standard CPR/congestion models typically violate (i). The slack-allocation rule is the unique linear proportional member that satisfies both while remaining boundary-continuous. This “No-Sucker-Loss” guarantee isolates cooperative agents from externalities created by over-claimants, a fairness property uncommon in standard CPR models.
Remark 1 (Why the No-Sucker-Loss guarantee is atypical).
For comparison with classic rules, write $C_{\mathrm{tot}}:=\sum_{j}C_{j}$ for total claims (distinct from our $X=\sum_{j}v_{j}$, total overage). In those models, even “cooperative” agents (those with $C_{j}\leq L_{j}$) can see their payoffs reduced when the system is under stress.
(i) Proportional rule on claims [4, 2]. When $C_{\mathrm{tot}}>I$, each agent receives a fraction $\lambda=I/C_{\mathrm{tot}}<1$ of their claim. Thus a cooperative agent with $C_{j}\leq L_{j}$ receives $\lambda C_{j}<C_{j}$, violating NLS.
(ii) Constrained equal awards (CEA) [5, 2]. Awards are $a_{j}=\min\{C_{j},\lambda\}$ with $\lambda$ chosen to exhaust the estate. With $C=(1,100,100)$ and $I=2$, $\lambda=2/3$ and the cooperative agent gets $a_{1}=2/3<1$, violating NLS.
(iii) Network proportional fairness [7, 8]. Allocations are jointly determined by a coupled optimization; holding $C_{j}$ fixed, increasing other users’ demands can strictly decrease agent $j$’s allocation, so there is no analogue of NLS.
In contrast, overage-proportional rationing applies reductions only to overages $(C_{j}-L_{j})_{+}$. Any agent with $C_{j}\leq L_{j}$ receives exactly $C_{j}$ regardless of others’ claims. These canonical families all fail NLS, indicating that the property is non-generic among standard rationing and congestion models.
3 Main properties
Proposition 2 (Cooperative frontier reproduced).
If all defect, then $I=0$, $\widehat{v}_{j}=0$, and $\pi_{j}=L_{j}$ for all $j$. If all cooperate, then $X=0$, $\pi_{j}=C_{j}$, and at $C=L$ the payoff vector is $(L_{j})_{j}$.
Proposition 3 (Dominant-strategy defection under bounds).
With $C_{j}\in[0,M]$ and the linear rule ($\alpha=1$), for any fixed $C_{-j}$ the map $C_{j}\mapsto\pi_{j}(C_{j},C_{-j})$ is nondecreasing; thus $C_{j}^{\star}=M$ is a best reply independent of $C_{-j}$. (See Appendix C.)
Theorem 4 (Uniqueness of boundary continuity).
Within the $\alpha$-power family for the slack allocation mechanism, continuity at $X=I$ for all positive overage vectors holds iff $\alpha=1$. (Appendix A.)
Theorem 5 (DS is Strong Nash under TU, therefore CPNE).
At the dominant-strategy outcome (all defect), under transferable utility within coalitions, no coalition $K$ can achieve a strict Pareto improvement by deviating; hence the profile is a Strong Nash equilibrium. By Bernheim et al. [1], Strong Nash implies Coalition-Proof Nash Equilibrium.
Proof.
Let $C^{\mathrm{DS}}=(M,\dots,M)$ with $C_{j}\in[0,M]$ (Prop. 3). Then $I=0$ and by Section 2 we have $\pi_{j}(C^{\mathrm{DS}})=L_{j}$ for all $j$, hence for any coalition $K\subseteq N$,
$\sum_{i\in K}\pi_{i}(C^{\mathrm{DS}})=\sum_{i\in K}L_{i}.$
Fix any coalition $K$ and any deviation $C^{\prime}_{K}$. The post-deviation profile is $C^{\prime}=(C^{\prime}_{K},\,C^{\mathrm{DS}}_{-K})$. Since the complement K-K continues to defect (plays $C^{\mathrm{DS}}_{-K}$), it generates no slack; therefore at $C^{\prime}$ we have $I=I_{K}$.
By Appendix B (the Case 2 argument applied with a defecting complement), whenever $I=I_{K}$ we have the coalition payoff bound
$\sum_{i\in K}\pi_{i}(C^{\prime})\;\leq\;\sum_{i\in K}L_{i}\;=\;\sum_{i\in K}\pi_{i}(C^{\mathrm{DS}}).$
Under transferable utility, a coalition deviation can make all its members weakly better and at least one strictly better only if its total payoff strictly increases. The bound shows this is impossible from $C^{\mathrm{DS}}$. Hence $C^{\mathrm{DS}}$ is a Strong Nash equilibrium under TU. Since every Strong Nash equilibrium is coalition-proof [1], $C^{\mathrm{DS}}$ is also a CPNE. ∎
Remark 6 (Equilibrium multiplicity under weak dominance).
The maximal-claiming profile $C^{\mathrm{DS}}=(M,\dots,M)$ is a dominant-strategy equilibrium. Because dominance is weak, other Nash equilibria exist. In particular, the cooperative profile $C=L$ is a Nash equilibrium: for any $j$ and any $C_{j}^{\prime}\geq L_{j}$, the induced $I^{\prime}=0$ yields $\pi_{j}(C_{j}^{\prime},C_{-j})=L_{j}=\pi_{j}(L_{j},C_{-j})$, so unilateral deviations are not profitable. Our welfare and coalition-proofness results are stated for the dominant-strategy outcome.
Remark 7 (Transferable utility (TU)).
We use TU in the standard sense: coalition members can make budget-balanced side-payments among themselves, so a deviation is evaluated by the coalition’s total payoff. Formally, for $K\subseteq N$ a deviation from $C$ to $C^{\prime}$ is feasible under TU iff there exist transfers $(t_{i})_{i\in K}$ with $\sum_{i\in K}t_{i}=0$ such that $\pi_{i}(C^{\prime})+t_{i}\geq\pi_{i}(C)$ for all $i\in K$, with strict inequality for at least one member.
4 Policy Implementation: Cap-and-Share with overage clearing
Interpret $L_{j}$ as allowances under cap-and-share (we index by $t$ only in this section and Appendix D), and take the linear rule ($\alpha=1$). In applications, the bound $M$ can represent a physical capacity, a regulatory limit, or a credit constraint; none of the results use more than $M>\max_{j}L_{j}$ (Appendix C).
A forward market clears expected buy/sell orders; at period end, realized emissions induce $(v_{j},s_{j})$ and clearing (1). Residual overage $(v_{j}-\widehat{v}_{j})_{+}$ is priced by a penalty collar $\kappa_{t}\in[\underline{\kappa},\overline{\kappa}]$ modulated by an endogenous scarcity factor $\Lambda_{t}$ (defined below). Appendix D (Prop. 10) shows that the mechanism is compatible with forward trading without perverse incentives.
Penalty and scarcity.
For period $t$, define the scarcity factor by
$\Lambda_{t}\;:=\;\begin{cases}0,&X^{t}=0,\\[3.0pt] \max\{0,(X^{t}-I^{t})/X^{t}\},&X^{t}>0,\end{cases}\qquad\Lambda_{t}\in[0,1].$
The per-unit penalty $\kappa_{t}\in[\underline{\kappa},\overline{\kappa}]$ is regulator-set (exogenous), while $\Lambda_{t}$ is endogenous (determined by realized $(X^{t},I^{t})$). If $X^{t}\leq I^{t}$, residual overage is fully covered (zero penalty); if $X^{t}>I^{t}$, a marginal unit of overage faces at least $\kappa_{t}\,\Lambda_{t}$ in penalty at clearing (Appendix D).
Risk & Governance.
The authority should emphasize robustness and auditability rather than discretion: (i) stress-test reporting and clearing against strategic misreporting and timing manipulation; (ii) publish, ex ante, the collar calibration and adjustment protocol (data sources and decision rules); (iii) monitor realized $(X^{t},I^{t},\Lambda_{t})$ against stated tolerances with review triggers for threshold breaches; and (iv) reserve a narrowly circumscribed emergency suspension rule that preserves budget balance and does not create profitable anticipatory deviations.
Appendix A. Overage-power family and boundary continuity
We generalize the slack-allocation mechanism by introducing an exponent $\alpha>0$ on overage shares. Players choose claims $C_{j}\geq 0$ against entitlements $L_{j}>0$ and we set, as in the main text,
$v_{j}=(C_{j}-L_{j})_{+},\qquad s_{j}=(L_{j}-C_{j})_{+},\qquad X=\sum_{m}v_{m},\qquad I=\sum_{m}s_{m}.$
Given a profile $C$, the clearing rule with exponent $\alpha$ allocates cooperative surplus $I$ to defectors ($v_{j}>0$) via
$\widehat{v}_{j}^{\alpha}(C)=\begin{cases}v_{j},&X\leq I,\\[6.0pt] \dfrac{I\,v_{j}^{\alpha}}{\sum_{m:v_{m}>0}v_{m}^{\alpha}},&X>I,\end{cases}$ (3)
and cooperators ($C_{j}\leq L_{j}$) receive their claim while defectors receive entitlement plus covered overage:
$\pi_{j}^{\alpha}(C)=\begin{cases}C_{j},&C_{j}\leq L_{j},\\[3.0pt] L_{j}+\widehat{v}_{j}^{\alpha}(C),&C_{j}>L_{j}.\end{cases}$
Boundary convention at $X=I$.
At the boundary we adopt the $X<I$ branch, i.e., $\widehat{v}_{j}^{\alpha}=v_{j}$. This preserves feasibility $\sum_{j:C_{j}>L_{j}}\widehat{v}_{j}^{\alpha}=I$ and budget balance. For $\alpha\neq 1$ the allocation generally differs from the $X>I$ branch, so a discontinuity remains; for $\alpha=1$ the two branches coincide and the rule is continuous.
Aggregate identity.
Let $\Pi^{\alpha}(C):=\sum_{j}\pi_{j}^{\alpha}(C)$. Since $\sum_{j:C_{j}\leq L_{j}}C_{j}=\sum_{j}\min\{C_{j},L_{j}\}$ and $\sum_{j:C_{j}>L_{j}}\widehat{v}_{j}^{\alpha}=\min\{X,I\}$, we have
$\Pi^{\alpha}(C)=\sum_{j}L_{j}-\max\{I-X,0\}.$ (4)
Hence in the scarcity region $X\geq I$ we have $\Pi^{\alpha}(C)=\sum_{j}L_{j}$, and in the slack region $X<I$ we have $\Pi^{\alpha}(C)=\sum_{j}L_{j}-(I-X)$. This coincides with the main-text budget identity (2).
Monotonicity in own claim under bounds.
Fix $C_{-j}$ and restrict $C_{j}\in[0,M]$. Set $y:=(C_{j}-L_{j})_{+}\in[0,M-L_{j}]$, $X=X_{-j}+y$, and $I=I_{-j}$; define $S_{-j}:=\sum_{m\neq j}v_{m}^{\alpha}$ (note: $S_{-j}\neq X_{-j}^{\alpha}$ in general). Then
$\pi_{j}^{\alpha}(y)=\begin{cases}L_{j}+y,&X_{-j}+y\leq I_{-j},\\[8.0pt] L_{j}+\dfrac{I_{-j}\,y^{\alpha}}{S_{-j}+y^{\alpha}},&X_{-j}+y>I_{-j},\end{cases}$
with the boundary value at $X_{-j}+y=I_{-j}$ set by the $X<I$ branch (boundary convention above).
On $0,(I_{-j}-X_{-j})_{+}]$ we have $\pi_{j}^{\alpha}(y)=L_{j}+y$ which is strictly increasing. On $((I_{-j}-X_{-j})_{+},\;M-L_{j}]$ we have
$\phi(y):=L_{j}+\frac{I_{-j}\,y^{\alpha}}{S_{-j}+y^{\alpha}},\qquad\phi^{\prime}(y)=\frac{\alpha\,I_{-j}\,S_{-j}\,y^{\alpha-1}}{(S_{-j}+y^{\alpha})^{2}}\;\geq 0,$
so $\pi_{j}^{\alpha}$ is nondecreasing there. For $\alpha=1$, the branches coincide at the boundary, hence $\pi_{j}^{1}(\cdot,C_{-j})$ is nondecreasing on $[0,M]$. In particular a best reply is attained at $C_{j}^{\star}=M$ for $\alpha=1$.
Theorem 8 (Uniqueness of boundary continuity).
For (3), the payoff map $C\mapsto\pi^{\alpha}(C)$ is continuous at profiles with $X=I$ for all positive overage vectors if and only if $\alpha=1$.
Proof.
Continuity away from $X=I$ is immediate. At $X=I$ with positive $v$, approaching from $X<I$ gives $\widehat{v}_{j}^{\alpha}=v_{j}$ while from $X>I$ gives $\widehat{v}_{j}^{\alpha}=I\,v_{j}^{\alpha}/\sum_{\ell}v_{\ell}^{\alpha}$. Equality for all positive $v$ forces $\frac{v_{j}^{\alpha}}{\sum_{\ell}v_{\ell}^{\alpha}}=\frac{v_{j}}{\sum_{\ell}v_{\ell}}$, which holds iff $\alpha=1$; conversely, for $\alpha=1$ the branches coincide. The boundary convention does not affect this characterization. ∎
Remark 9 (Uniqueness of the linear rule at the boundary).
On the scarcity side $X>I$, consider the $\alpha$-weighted alternative that rations overages by weights proportional to $v_{j}^{\alpha}$:
$\pi_{j}^{(\alpha)}\;=\;L_{j}\;+\;I\,\frac{v_{j}^{\alpha}}{\sum_{i}v_{i}^{\alpha}}.$
On the slack side $X<I$, payoffs are $\pi_{j}=L_{j}+v_{j}$. At the boundary $X=I$ the one-sided limits match for all overage vectors $v$ if and only if $\alpha=1$. For any $\alpha\neq 1$, equality fails except on the measure-zero set where all positive $v_{j}$ are equal, so a jump occurs at $X=I$.
Appendix B. Coalition-proofness at the DS outcome
We prove coalition-proofness for the linear rule ($\alpha=1$) under slack allocation.
Proof.
Let $K$ be any coalition and assume the complement K-K defects. We show that the coalition’s aggregate payoff at any deviation $C_{K}$ cannot exceed $\sum_{i\in K}L_{i}$.
Let $S:=\{\,i:C_{i}\leq L_{i}\,\}$ (cooperators) and $D:=\{\,i:C_{i}>L_{i}\,\}$ (defectors). Write $X_{K}=\sum_{j\in K}v_{j}$, $_{-K}=\sum_{j\notin K}v_{j}, X=X$$X_{K}+X_{-K}$, and $I_{K}=\sum_{i\in K}s_{i}$. Since K-K defects, their slack is 0, so $I=I_{K}$.
The coalition’s aggregate payoff is derived as follows. Note that for $i\in S$, $C_{i}=L_{i}-s_{i}$.
$\sum_{i\in K}\pi_{i}$ $=\sum_{i\in K\cap S}C_{i}\;+\;\sum_{j\in K\cap D}\big{(}L_{j}+\widehat{v}_{j}\big{)}$
$=\sum_{i\in K\cap S}(L_{i}-s_{i})\;+\;\sum_{j\in K\cap D}L_{j}\;+\;\sum_{j\in K\cap D}\widehat{v}_{j}$
$=\sum_{i\in K}L_{i}\;-\;\sum_{i\in K\cap S}s_{i}\;+\;\sum_{j\in K\cap D}\widehat{v}_{j}.$
Since $s_{i}=0$ for $i\in D$, we have $\sum_{i\in K\cap S}s_{i}=\sum_{i\in K}s_{i}=I_{K}$. Thus,
$\sum_{i\in K}\pi_{i}=\sum_{i\in K}L_{i}\;-\;I_{K}\;+\;\sum_{j\in K\cap D}\widehat{v}_{j}.$ (5)
We bound the last term in the two regions.
Case 1: $X\leq I$. All overage is covered, so $\sum_{j\in K\cap D}\widehat{v}_{j}=X_{K}$. Then by (5), $\sum_{i\in K}\pi_{i}=\sum_{i\in K}L_{i}+(X_{K}-I_{K}).$ But $X\leq I$ and $I=I_{K}$ imply $X_{K}\leq I_{K}-X_{-K}\leq I_{K}$, hence $X_{K}-I_{K}\leq-X_{-K}\leq 0$. Therefore $\sum_{i\in K}\pi_{i}\leq\sum_{i\in K}L_{i}$.
Case 2: $X>I$. Coverage is proportional: $\sum_{j\in K\cap D}\widehat{v}_{j}=I\cdot\frac{X_{K}}{X}=I_{K}\cdot\frac{X_{K}}{X}$. Thus by (5),
$\sum_{i\in K}\pi_{i}$ $=\sum_{i\in K}L_{i}-I_{K}+I_{K}\tfrac{X_{K}}{X}$
$=\sum_{i\in K}L_{i}-I_{K}\big{(}1-\tfrac{X_{K}}{X}\big{)}$
$=\sum_{i\in K}L_{i}-I_{K}\tfrac{X_{-K}}{X}\;\leq\;\sum_{i\in K}L_{i}.\qed$
Appendix C. Bounded-action regularization
Fix $M>0$. Each player $j$ chooses $C_{j}\in[0,M]$.
Assumption (large action bound). Throughout Appendix C and any results that invoke it, take $M>\max_{j}L_{j}$, so that the maximal claim $M$ constitutes defection; if this fails, replace “defection” with “maximal claim” in the statements without altering the analysis.
Let $v_{j}=(C_{j}-L_{j})_{+}$, $s_{j}=(L_{j}-C_{j})_{+}$, $X=\sum_{m}v_{m}$, $I=\sum_{m}s_{m}$. For $\alpha=1$, $\widehat{v}_{j}=v_{j}$ if $X\leq I$ and $\widehat{v}_{j}=(I/X)v_{j}$ if $X>I$. Payoffs are $\pi_{j}=C_{j}$ when $C_{j}\leq L_{j}$, and $\pi_{j}=L_{j}+\widehat{v}_{j}$ when $C_{j}>L_{j}$.
Proposition C. (i) Best replies exist. (ii) For any fixed $C_{-j}$, $\pi_{j}(C_{j},C_{-j})$ is nondecreasing on $[0,M]$, hence a best reply is $C_{j}^{\star}=M$.
Proof.
Fix $C_{-j}$. We first establish continuity. On $0,L_{j}]$, $\pi_{j}(C_{j},C_{-j})=C_{j}$. For the region $C_{j}\geq L_{j}$, write $y:=(C_{j}-L_{j})_{+}\in[0,M-L_{j}]$, $X=X_{-j}+y$, $I=I_{-j}$. Then
$\pi_{j}(y)=\begin{cases}L_{j}+y,&X_{-j}+y\leq I_{-j},\\[4.0pt] L_{j}+\dfrac{I_{-j}y}{X_{-j}+y},&X_{-j}+y>I_{-j}.\end{cases}$
We check continuity at the switching point $y^{\star}=(I_{-j}-X_{-j})_{+}$. If $y^{\star}>0$, then $y^{\star}=I_{-j}-X_{-j}$, which implies $X_{-j}+y^{\star}=I_{-j}$. The two branches evaluate to:
Branch 1 (slack): $\quad L_{j}+y^{\star},$
Branch 2 (scarcity): $\quad L_{j}+\frac{I_{-j}y^{\star}}{X_{-j}+y^{\star}}=L_{j}+\frac{I_{-j}y^{\star}}{I_{-j}}=L_{j}+y^{\star}.$
If $y^{\star}=0$ (i.e., $I_{-j}\leq X_{-j}$), the payoff function near $y=0$ is continuous, and both expressions evaluate to $L_{j}$ at $y=0$. Thus the branches agree at $y^{\star}$, and $\pi_{j}(\cdot,C_{-j})$ is continuous on $[0,M]$.
(i) Since $\pi_{j}(\cdot,C_{-j})$ is continuous on the compact interval $[0,M]$, a maximizer exists by the Weierstrass extreme value theorem.
(ii) We establish monotonicity. On $0,L_{j}]$, $\pi_{j}$ is strictly increasing. For $C_{j}>L_{j}$: On $0,y^{\star}]$, $\pi_{j}(y)=L_{j}+y$ is strictly increasing. On $(y^{\star},M-L_{j}]$, $\frac{\mathrm{d}}{\mathrm{d}y}\!\left(L_{j}+\frac{I_{-j}y}{X_{-j}+y}\right)=\frac{I_{-j}X_{-j}}{(X_{-j}+y)^{2}}\geq 0$, so $\pi_{j}$ is nondecreasing. Hence $\pi_{j}$ is nondecreasing on $[0,M]$ and a best reply is $C_{j}^{\star}=M$. ∎
Appendix D. Penalty-Collar Economics: No Gain from Strategic Over-Emission
By a penalty collar we mean a regulated interval $\underline{\kappa},\overline{\kappa}]$ for the per-unit penalty applied to uncovered residual overage at clearing; the realized period-$t$ penalty is $\kappa_{t}\in[\underline{\kappa},\overline{\kappa}]$.
Setup and notation.
Fix period $t$. Each entity $i$ has entitlement $L_{i}^{t}>0$ and realizes usage (claims) $C_{i}^{t}\geq 0$. Define overage $v_{i}:=(C_{i}^{t}-L_{i}^{t})_{+}$, slack $s_{i}:=(L_{i}^{t}-C_{i}^{t})_{+}$, aggregate $X^{t}=\sum_{i}v_{i}$, $I^{t}=\sum_{i}s_{i}$. For any $j$, let $V_{-j}:=\sum_{m\neq j}v_{m}$ denote the aggregate overage of others, so $X^{t}=v_{j}+V_{-j}$.
The scarcity factor $\Lambda_{t}$ is defined in Section 4. End-of-period clearing covers defectors’ overage proportionally: $\widehat{v}_{i}=v_{i}$ if $X^{t}\leq I^{t}$, else $\widehat{v}_{i}=(I^{t}/X^{t})\,v_{i}$. Residual overage is $r_{i}:=(v_{i}-\widehat{v}_{i})_{+}$. Note that $r_{i}=\Lambda_{t}v_{i}$ when $X^{t}>I^{t}$ and 0 otherwise.
Let $p_{\tau}$ be the forward price at decision time $\tau<t$, conditional on information $\mathcal{F}_{\tau}$.
Prices and calibration parameters.
All expectations below are conditional on the information set defined in the next paragraph. Let $p_{t}$ denote the period-$t$ spot price at clearing, and let $\overline{p}_{t}$ be a publicly announced upper bound on $p_{t}$ (e.g., an auction reserve or penalty ceiling). Assumption (expected scarcity). There exists $\underline{\lambda}\in(0,1]$ such that $\mathbb{E}[\Lambda_{t}\mid\mathcal{F}_{\tau}]\geq\underline{\lambda}$. This assumption is used only in the collar-calibration corollary below.
Information set.
For period $t$, let $\mathcal{F}_{\tau}$ denote the public information available by decision time $\tau<t$: (i) entitlements $\{L_{i}^{t}\}$; (ii) policy parameters $(\underline{\kappa},\overline{\kappa})$; (iii) forward orders/positions and any other public signals observed by $\tau$ that bear on the period-$t$ aggregates $(X^{t},I^{t})$ and on the realized penalty $\kappa_{t}$. Expectations $\mathbb{E}[\cdot\mid\mathcal{F}_{\tau}]$ are conditional on that information. At clearing, $\Lambda_{t}$ and $\kappa_{t}$ are $\mathcal{F}_{t}$-measurable (but need not be $\mathcal{F}_{\tau}$-measurable for $\tau<t$).
Proposition 10 (Expected marginal cost of waiting).
For any defector $j$ with overage $v_{j}$ at time $t$, the expected unit cost of creating one additional unit by waiting for clearing is at least
$\mathbb{E}\!\left[\kappa_{t}\,\Lambda_{t}\,\middle|\,\mathcal{F}_{\tau}\right].$
Proof.
Write $P_{j}(v_{j}):=\kappa_{t}\,r_{j}(v_{j})$ for $j$’s penalty at clearing, where $r_{j}=(v_{j}-\widehat{v}_{j})_{+}$. The expected unit cost of creating one more unit by waiting equals the conditional expectation of the right marginal $\partial P_{j}/\partial v_{j}$ holding $(V_{-j},I^{t})$ fixed.
If $X^{t}\leq I^{t}$, then $\Lambda_{t}=0$ in a neighborhood and $r_{j}\equiv 0$, so $\frac{\partial r_{j}}{\partial v_{j}}=0=\Lambda_{t}$, hence $\frac{\partial P_{j}}{\partial v_{j}}=\kappa_{t}\,\frac{\partial r_{j}}{\partial v_{j}}\geq\kappa_{t}\,\Lambda_{t}$.
If $X^{t}>I^{t}$, then $\Lambda_{t}=\frac{(X^{t}-I^{t})}{X^{t}}=1-\frac{I^{t}}{X^{t}}$ with $X^{t}=v_{j}+V_{-j}$. Differentiating w.r.t. $v_{j}$ gives
$\frac{\partial\Lambda_{t}}{\partial v_{j}}=\frac{I^{t}}{(X^{t})^{2}}\;\geq\;0.$
Since $r_{j}=\Lambda_{t}\,v_{j}$ in this region,
$\frac{\partial r_{j}}{\partial v_{j}}=\Lambda_{t}+v_{j}\,\frac{\partial\Lambda_{t}}{\partial v_{j}}\;\geq\;\Lambda_{t},$
and therefore $\frac{\partial P_{j}}{\partial v_{j}}=\kappa_{t}\,\frac{\partial r_{j}}{\partial v_{j}}\geq\kappa_{t}\,\Lambda_{t}$.
At the kink $X^{t}=I^{t}$ the right derivative exists and the same inequality holds by the above cases. Taking conditional expectations yields
$\mathbb{E}\!\left[\frac{\partial P_{j}}{\partial v_{j}}\,\middle|\,\mathcal{F}_{\tau}\right]\;\geq\;\mathbb{E}\!\left[\kappa_{t}\,\Lambda_{t}\,\middle|\,\mathcal{F}_{\tau}\right],$
which proves the claim. ∎
Lemma 11 (No benefit from inflating overage).
Fix $v_{-j}$ and $I^{t}$ with $X^{t}>I^{t}$. With $X^{t}=v_{j}+V_{-j}$ and $\widehat{v}_{j}=I^{t}\,v_{j}/X^{t}$,
$\frac{\mathrm{d}}{\mathrm{d}v_{j}}r_{j}(v_{j})=1-\frac{\mathrm{d}}{\mathrm{d}v_{j}}\Big{(}\frac{I^{t}v_{j}}{X^{t}}\Big{)}=1-I^{t}\frac{X^{t}-v_{j}}{(X^{t})^{2}}\geq 1-\frac{I^{t}}{X^{t}}=\Lambda_{t},$
holding $I^{t},V_{-j}$ fixed. Thus $r_{j}$ is strictly increasing and the incremental penalty is at least $\kappa_{t}\,\Lambda_{t}$ per unit.
Corollary 12 (Collar calibration kills “wait-and-emit” arbitrage).
If the authority sets $\kappa_{t}\geq\overline{p}_{t}$ (auction reserve or price cap) and publishes $\mathbb{E}[\Lambda_{t}\mid\mathcal{F}_{\tau}]\geq\underline{\lambda}>0$, then $\mathbb{E}[\kappa_{t}\,\Lambda_{t}\mid\mathcal{F}_{\tau}]\geq\overline{p}_{t}\,\underline{\lambda}$. If $\overline{p}_{t}\,\underline{\lambda}\geq p_{\tau}$, forward purchase is weakly cheaper in expectation than waiting; strict if >>.
Budget balance reminder.
When $X^{t}\geq I^{t}$, the clearing is budget balanced, i.e., $\sum_{i}\pi_{i}^{t}=\sum_{i}L_{i}^{t}$ (cf. Eq. (2); here $\pi_{i}^{t}$ denotes the period-$t$ payoff.